# Offline and Online Use of Interval and Set-Based Approaches for Control and State Estimation: A Selection of Methodological Approaches and Their Application

### v3

@@ 4.1. Goal of the Control Synthesis
The main drawback of neglecting this information about uncertainty is the lack of a robustness measure of the resulting control. This neglection may lead to scenarios where the determined control signal does only stabilize the system model for parts of the possible parameter ranges, while the resulting dynamics are unstable for others. This is especially critical for open-loop unstable system models such as the inverted pendulum system in Figure 3, for which a model predictive controller was designed in [30]. This controller simultaneously aims at a swing-up of the pendulum arm $S_{2}$ by rotation of the arm $S_{1}$ and holding $S_{2}$ it in its unstable upright equilibrium by means of pure inertia coupling.
@@ 4.3. Interval-Based Nonlinear Model Predictive Control
  • Filtering and branching: This first step provides a sequence of guaranteed input interval boxes at each time-step $t_{k}$ over the prediction horizon $N_{\mathrm{p}}$, denoted as $[\mathbf{U}(t_{k})]=[\mathbf{{u}}(t_{k})]\times[\mathbf{{u}}{(t_{k+1})}]\times\ldots\times[\mathbf{{u}}(t_{k+N_{\mathrm{p}}-1})]$. Branching and filtering procedures allow the computation of safe input intervals [^7] For the implementation of this branching algorithm we assume piecewise constant input intervals. Therefore, the more general notation $[\mathbf{{u}}(t_{k})]$ is subsequently replaced by $[\mathbf{{u}}_{k}]$ to indicate control signals only changing their value at the points in time $t_{k}$. along the receding time horizon. In this context, safety means that the state constraints that are specified in the form of admissible interval in each time step (i.e., $\forall j,\left[\mathbf{{x}}(t_{j})\right]\subseteq[\mathbf{x}_{\min}(t_{j}),\mathbf{x}_{\max}(t_{j})]$ are guaranteed to be obeyed despite the parameter uncertainty, where $\mathbf{x}_{\min}(t_{j})$ and $\mathbf{x}_{\max}(t_{j})$ are the bounds for the admissible domain for each state variable). In addition, the desired convergence to the reference interval (i.e., $[\mathbf{x}_{j}]\rightarrow[\mathbf{x}_{\mathrm{r}}]$) needs to be ensured.
@@ 4.4. Formulation and Minimization of Interval Cost Functions
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}F(\mathbf{x}(t),\mathbf{u}(\tau))=\left(\mathbf{x}(t)-\mathbf{x}_{\mathrm{r}}\right)^{T}\mathbf{Q}\left(\mathbf{x}(t)-\mathbf{x}_{\mathrm{r}}\right)+\mathbf{u}(t)^{T}\mathbf{R}{\mathbf{u}}(t)\enspace,}$ (57)
@@
Figure 4 gives an overview of a possible extension of the interval-based nonlinear model predictive control procedure. This extension is represented by a robust pre-stabilization of open-loop unstable plants with the help of linear matrix inequality (LMI) techniques. In such cases, the interval-based optimization procedure does not directly provide the overall control sequence but rather a feedforward control sequence $[\mathbf{u}_{\mathrm{ff}}]$ that allows for optimizing the tracking behavior within the operating domain in which the underlying feedback controller possesses provable asymptotic stability properties [30]. For an application of this control procedure to the swing-up control of a rotary inverse pendulum sketched in Sec. 4.1, the reader is referred to [30].
@@ 6.3. Ellipsoidal Calculus for the Guaranteed Preduction of Covariance Ellipsoids
$\mathbf{b}_{k}=\left({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{\Phi}\left({\mathbf{z}_{k},\mathbf{p}}\right)}-\tilde{\boldsymbol{\Phi}}}\right)\cdot\boldsymbol{\mu}_{k}\in\left[{\mathbf{b}_{k}}\right]$ (84)
@@ 6.4.1. Prediction Step
in which joint distributions over the two subsequent trial are denoted by the superscript $i|i+1$, as well as of the trial-independent noise covariance matrix $\mathbf{C}_{\mathbf{w},k}$, and the magnification factor $r\geq 1$ as a user-defined degree of freedom. Here, in analogy to the Kalman Filter, the term innovation step refers to the operation in which the predicted state information obtained from an open-loop evaluation of the system model into the future is corrected by taking into account measured data.
@@ 6.4.2. Innovation Step
$\mathbf{C}_{k}^{\mathrm{p},i|i+1}=\begin{bmatrix}\mathbf{C}_{\mathrm{A},k}^{\mathrm{p}}&\mathbf{C}_{\mathrm{B},k}^{\mathrm{p}}\\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left({\mathbf{C}_{\mathrm{B},k}^{\mathrm{p}}}\right)^{T}}&\mathbf{C}_{\mathrm{C},k}^{\mathrm{p}}\end{bmatrix}\enspace.$ (106)
@@
$\mathbf{P}\cdot{\tilde{\boldsymbol{\Gamma}}_{k}^{\mathrm{e}}\cdot\left({\tilde{\boldsymbol{\Gamma}}_{k}^{\mathrm{e}}}\right)^{T}}\cdot\mathbf{P}^{T}+\tilde{\mathbf{C}}_{\mathbf{v},k}=\begin{bmatrix}\mathbf{C}_{\mathrm{A},k}&\mathbf{C}_{\mathrm{B},k}\\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{C}_{\mathrm{B},k}^{T}}&\mathbf{C}_{\mathrm{C},k}\end{bmatrix}$ (107)
@@
It should be noted at this point that setting all coupling terms between the iterations $i$ and $i+1$ leads exactly to the same gain matrix that would be obtained when designing a Kalman filter if all matrices in (90) are state-independent and only depend on precisely known parameters.

### v4

@@ 4.1. Goal of the Control Synthesis
The main drawback of neglecting this information about uncertainty is the lack of a robustness measure of the resulting control. This neglection may lead to scenarios where the determined control signal does only stabilize the system model for parts of the possible parameter ranges, while the resulting dynamics are unstable for others. This is especially critical for open-loop unstable system models such as the inverted pendulum system in Figure 3, for which a model predictive controller was designed in [30]. This controller simultaneously aims at a swing-up of the pendulum arm $S_{2}$ by rotation of the arm $S_{1}$ and holding $S_{2}$ in its unstable upright equilibrium by means of pure inertia coupling.
@@ 4.3. Interval-Based Nonlinear Model Predictive Control
  • Filtering and branching: This first step provides a sequence of guaranteed input interval boxes at each time-step $t_{k}$ over the prediction horizon $N_{\mathrm{p}}$, denoted as $[\mathbf{U}(t_{k})]=[\mathbf{{u}}(t_{k})]\times[\mathbf{{u}}{(t_{k+1})}]\times\ldots\times[\mathbf{{u}}(t_{k+N_{\mathrm{p}}-1})]$. Branching and filtering procedures allow the computation of safe input intervals [^7] For the implementation of this branching algorithm we assume piecewise constant input intervals. Therefore, the more general notation $[\mathbf{{u}}(t_{k})]$ is subsequently replaced by $[\mathbf{{u}}_{k}]$ to indicate control signals only changing their value at the points in time $t_{k}$. along the receding time horizon. In this context, safety means that the state constraints that are specified in the form of admissible intervals in each time step (i.e., $\forall j,\left[\mathbf{{x}}(t_{j})\right]\subseteq[\mathbf{x}_{\min}(t_{j}),\mathbf{x}_{\max}(t_{j})]$ are guaranteed to be obeyed despite the parameter uncertainty, where $\mathbf{x}_{\min}(t_{j})$ and $\mathbf{x}_{\max}(t_{j})$ are the bounds for the admissible domain for each state variable). In addition, the desired convergence to the reference interval (i.e., $[\mathbf{x}_{j}]\rightarrow[\mathbf{x}_{\mathrm{r}}]$) needs to be ensured.
@@ 4.4. Formulation and Minimization of Interval Cost Functions
${F(\mathbf{x}(t),\mathbf{u}(\tau))=\left(\mathbf{x}(t)-\mathbf{x}_{\mathrm{r}}\right)^{T}\mathbf{Q}\left(\mathbf{x}(t)-\mathbf{x}_{\mathrm{r}}\right)+\mathbf{u}(t)^{T}\mathbf{R}{\mathbf{u}}(t)\enspace,}$ (57)
@@
Figure 4 gives an overview of a possible extension of the interval-based nonlinear model predictive control procedure. This extension is represented by a robust pre-stabilization of open-loop unstable plants with the help of linear matrix inequality (LMI) techniques. In such cases, the interval-based optimization procedure does not directly provide the overall control sequence but rather a feedforward control action $[\mathbf{u}_{\mathrm{ff}}]$ that allows for optimizing the tracking behavior within the operating domain in which the underlying feedback controller possesses provable asymptotic stability properties [30]. For an application of this control procedure to the swing-up control of a rotary inverse pendulum sketched in Sec. 4.1, the reader is referred to [30].
@@ 6.3. Ellipsoidal Calculus for the Guaranteed Preduction of Covariance Ellipsoids
$\mathbf{b}_{k}=\left({{\boldsymbol{\Phi}\left({\mathbf{z}_{k},\mathbf{p}}\right)}-\tilde{\boldsymbol{\Phi}}}\right)\cdot\boldsymbol{\mu}_{k}\in\left[{\mathbf{b}_{k}}\right]$ (84)
@@ 6.4.1. Prediction Step
in which joint distributions over the two subsequent trials are denoted by the superscript $i|i+1$, as well as of the trial-independent noise covariance matrix $\mathbf{C}_{\mathbf{w},k}$, and the magnification factor $r\geq 1$ as a user-defined degree of freedom. Here, in analogy to the Kalman Filter, the term innovation step refers to the operation in which the predicted state information obtained from an open-loop evaluation of the system model into the future is corrected by taking into account measured data.
@@ 6.4.2. Innovation Step
$\mathbf{C}_{k}^{\mathrm{p},i|i+1}=\begin{bmatrix}\mathbf{C}_{\mathrm{A},k}^{\mathrm{p}}&\mathbf{C}_{\mathrm{B},k}^{\mathrm{p}}\\ {\left({\mathbf{C}_{\mathrm{B},k}^{\mathrm{p}}}\right)^{T}}&\mathbf{C}_{\mathrm{C},k}^{\mathrm{p}}\end{bmatrix}\enspace.$ (106)
@@
$\mathbf{P}\cdot{\tilde{\boldsymbol{\Gamma}}_{k}^{\mathrm{e}}\cdot\left({\tilde{\boldsymbol{\Gamma}}_{k}^{\mathrm{e}}}\right)^{T}}\cdot\mathbf{P}^{T}+\tilde{\mathbf{C}}_{\mathbf{v},k}=\begin{bmatrix}\mathbf{C}_{\mathrm{A},k}&\mathbf{C}_{\mathrm{B},k}\\ {\mathbf{C}_{\mathrm{B},k}^{T}}&\mathbf{C}_{\mathrm{C},k}\end{bmatrix}$ (107)
@@
It should be noted at this point that setting all coupling terms between the iterations $i$ and $i+1$ to zero leads exactly to the same gain matrix that would be obtained when designing a Kalman filter if all matrices in (90) are state-independent and only depend on precisely known parameters.